### v4

# Towards Stable Machine Learning Model Retraining via Slowly Varying Sequences

Abstract
We consider the task of retraining machine learning (ML) models when new batches of data become available. Existing methods focus largely on greedy approaches to find the best-performing model for each batch, without considering the stability of the model’s structure across retraining iterations. In this study, we propose a methodology for finding sequences of ML models that are stable across retraining iterations. We develop a mixed-integer optimization formulation that is guaranteed to recover Pareto optimal models (in terms of the predictive power-stability trade-off) and an efficient polynomial-time algorithm that performs well in practice. We focus on retaining consistent analytical insights — which is important to model interpretability, ease of implementation, and fostering trust with users — by using custom-defined distance metrics that can be directly incorporated into the optimization problem. Our method shows stronger stability than greedily trained models with a small, controllable sacrifice in predictive power, as evidenced through a real-world case study in a major hospital system in Connecticut.
1 Introduction
The adoption of ML models in high-stakes decision spaces such as healthcare [1] increases the need for interpretability and model understanding [2, 3, 4]. Expert users develop trust in AI-driven decision support systems based on both the systems’ alignment with the users’ personal expertise [5, 6] and the systems’ historical performance [7]. Importantly, trust is built over time and requires deep understanding of the underlying model and the resulting analytical insights [8]. In real-world production pipelines, ML models need to be frequently updated as more data becomes available [9]. Significant changes in the model’s structure and the resulting analytical insights [10] may lead to skepticism and hesitation for adoption. Especially in healthcare, such changes may challenge both physicians’ trust in AI systems and patients’ confidence in medical decision-making [11, 12], and may even lead to legal difficulties as a result of incidents of error [13].
Multiple avenues have been proposed in the literature to address challenges associated with model retraining, including continual learning [14, 15], online learning [16], and others [17, 18, 19, 20]. Existing methods focus largely on maintaining the predictive power of the retrained models in the face of, e.g., data distributional shifts. Such methods do not, however, ensure that the decision path of a tree-based prediction or the important features selected by a model remain stable across different retraining iterations, and therefore do not guarantee retention of analytical insights. In this work, we explicitly incorporate this requirement into the (re)training process.
Retraining (batch learning) differs from online learning: the former periodically retrains the model from scratch with each new data batch; the latter continuously/incrementally updates the model as new data arrives. Thus, a batch learning-based retraining strategy allows for a globally optimal solution, as retraining uses the entire database, whereas online learning, often to ensure faster updates, incrementally incorporates only a portion of the data, hence providing greedier solutions [21].
The problem of developing stable ML models is not new in the statistics, computer science, and operations research literature [22, 10, 23]. Various methodologies have recently been proposed to either train structurally stable, interpretable models [24, 25, 26, 27], or black-box models that produce stable feature interactions [28, 29]. In this paper, we develop a unifying, model-agnostic framework for training sequences of models that remain stable upon retraining with new data.
We summarize the main contributions of this paper: (i) We develop a general framework aimed at stabilizing ML model structures upon retraining with new data. We propose an exact mixed-integer optimization (MIO) formulation as well as a heuristic, restricted formulation for the problem. Using the latter, we develop a practical, model-agnostic algorithm for retraining ML models. We also provide theoretical guarantees on the Pareto optimality and generalization of the solutions obtained by our methodology. (ii) We demonstrate across a variety of models (including logistic regression, classification trees, and boosted trees) that our algorithm gives a significant boost in model stability with a small and controllable sacrifice of predictive power and computational efficiency compared to existing greedy approaches. (iii) We show our method’s applicability in an existing production pipeline used in a large US hospital for mortality prediction.
2 A Methodology for Retraining Machine Learning Models
2.1 Problem Setting
Consider an observational dataset accumulated over time composed of triplets $(\bm{x_{i}},y_{i},t_{i})_{i=1}^{N}$, where $\bm{x_{i}}\in\mathcal{X}\subseteq\mathbb{R}^{p}$ represents the feature vector, $y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$ is the target variable, and $t_{i}\in\mathbb{R}$ is the time at which data point $i$ was obtained. The dataset is divided into a series of batches at times $T_{1}<T_{2}<\ldots<T_{B}$. The dataset available at time $T_{b},\ b\in[B]:=\{1,\ldots,B\},$ is defined as $D_{b}=\{(x_{i},y_{i}):t_{i}\leq T_{b},i\in[N]\}.$ Thus, we have that $D_{1}\subseteq D_{2}\subseteq\ldots\subseteq D_{B}$ and $\bm{D}=(D_{1},\dots,D_{B})$. We assume that $B\geq 2$ and that, at model training time (denoted by $t$), all $B$ batches are available, i.e., $t\geq T_{B}$. In Section 2.5, we discuss how to adaptively retrain the model using batches obtained after time $t$. The goal is to find models $\bm{f}:=(f_{1},f_{2},\ldots,f_{B})\in\mathcal{F}\times...\times\mathcal{F}:=\mathcal{F}^{B}$ with $f_{b}:\mathcal{X}\mapsto\mathcal{Y}$ for $b\in[B]$ such that the following two objectives are satisfied:
Predictive power: The first objective is to ensure that each model $f_{b}$ in the sequence has small expected loss in terms of prediction quality. We denote by $\mathbb{P}(D_{b})$ the distribution of the training data in $D_{b}$ and $\ell:\mathcal{Y}\times\mathcal{Y}\mapsto\mathbb{R}_{\geq 0}$ the bounded loss function measuring, e.g., regression or classification accuracy; concretely, $\ell(y,f_{b}(\bm{x}))$ quantifies the quality of model $f_{b}$’s prediction on input $\bm{x}$ for $y$. Then, we seek for a sequence whereby each model (individually) achieves low expected loss $\mathcal{L}_{P}(f_{b},D_{b})=\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}(D_{b})}[\ell(y,f_{b}(\bm{x}))].$ The overall loss of the sequence is then denoted by $\mathcal{L}_{P}(\bm{f},\bm{D})=\sum_{b=1}^{B}\mathcal{L}_{P}(f_{b},D_{b}).$ In practice, we approximate the expected loss using the empirical loss ^$\hat{\mathcal{L}}_{P}(f_{b},D_{b})=\frac{1}{N_{b}}\sum_{i:t_{i}\leq T_{b}}[\ell(y_{i},f_{b}(\bm{x}_{i}))]$. The next assumption bounds the approximation quality as function of the sample size, the function class complexity, and the desired confidence:
Assumption 1.
The predictive power objective satisfies a generalization bound of the form: $\forall\delta\in(0,1)$, $\mathbb{P}\left(\forall f\in\mathcal{F}:|\mathcal{L}_{P}(f,D)-\hat{\mathcal{L}}_{P}(f,D)|\leq C(N,\mathcal{F},\delta)\right)\geq 1-\delta$.
Stability: The second objective is to ensure that consecutive models do not vary significantly, maintaining stability over time. Formally, we define a stability metric $d:\mathcal{F}\times\mathcal{F}\mapsto\mathbb{R}_{\geq 0}$ so that $d(f_{i},f_{j})$ measures the distance (change) between models $i$ and $j$. We remark that the stability metric could be based on parameters (in parametric models), structure (in tree-based models), or even predictions. Then, for each pair of consecutive models, we minimize $\mathcal{L}_{S}(\bm{f})=\sum_{b=1}^{B-1}d(f_{b},f_{b+1})$.
Observation 1.
The stability loss is data-independent, i.e., $\mathcal{L}_{S}(\bm{f},\bm{D})=\mathcal{L}_{S}(\bm{f})$ for all data instances $\bm{D}$, so the stability objective generalizes trivially, i.e., ^$\hat{\mathcal{L}}_{S}(\bm{f})=\mathcal{L}_{S}(\bm{f})$.
In practice, we are primarily interested in model $f_{B}$ as this model has been trained on the full training set and is the one that will be used in production. Nevertheless, we choose to train the full sequence $f_{1},\ldots,f_{B}$ of $B$ models since, by doing so, (i) we favor models that change smoothly upon the addition of new batches of data, and (ii) we are able to observe changing patterns, which is crucial for interpretability. Our approach thus relies on imposing stability by process.
2.2 A General Formulation
The objectives stated in Section 2.1 naturally give rise to a multi-objective optimization problem [30]: we search for a sequence $\bm{f}$ that “performs well” in terms of both ^$\hat{\mathcal{L}}_{P}$ and ^$\hat{\mathcal{L}}_{S}$. This is commonly achieved by minimizing (i) a weighted combination of the objectives, i.e., $\lambda\hat{\mathcal{L}}_{P}+(1-\lambda)\hat{\mathcal{L}}_{S}$, where the weight $\lambda\in[0,1]$ is selected based on the learner’s preferences (assuming these are known), or (ii) a preference-agnostic, min-max type formulation [31]. In our setting, we often want to explicitly control “the price of stability” with respect to predictive power. This is particularly relevant in healthcare applications, where the accuracy requirements for using a model in production are stringent. Therefore, we opt for an $\epsilon$-constrained formulation: given a suboptimality tolerance parameter $\alpha\geq 0$, the predictive error of each model $f_{b}$ should not be worse than the best model for batch $D_{b}$ by more than a multiplicative factor of $\alpha$. Recalling our assumption that, at training time, we have already received $B\geq 2$ batches of data (if this is not the case, we artificially split the training data into $B$ batches), we obtain the following formulation:
$\begin{split}\min_{f_{1},\ldots,f_{B}\in\mathcal{F}^{B}}&\quad\hat{\mathcal{L}}_{S}(\bm{f})\ :=\ \sum_{b=1}^{B-1}d(f_{b},f_{b+1})\\ \text{s.t.}&\quad\hat{\mathcal{L}}_{P}(f_{b},D_{b})\leq(1+\alpha)\cdot\min_{f\in\mathcal{F}}\hat{\mathcal{L}}_{P}(f,D_{b}):=\epsilon_{b}\quad\forall b\in[B].\end{split}$ (1)
To analyze the quality of solutions to Problem (1), we adapt the definition of Pareto optimality:
Definition 1.
A sequence $\bm{f}\in\mathcal{F}^{B}$ is weakly Pareto optimal (WPO) if there is no other sequence $\bm{f^{\prime}}$ such that ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})<\hat{\mathcal{L}}_{P}(f_{b},D_{b})\ \forall b\in[B]$ and ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})<\hat{\mathcal{L}}_{S}(\bm{f})$. $\bm{f}$ is Pareto optimal (PO) if the inequality is strict for at least one objective and non-strict for the remaining ones.
The next results show that any solution to Problem (1) is WPO, whereas a unique solution is PO:
Theorem 1.
A sequence $\bm{f^{*}}$ obtained by solving Problem (1) is WPO.
Proof.
(Proof of Theorem 1) Let $\bm{f^{*}}$ be a solution to Problem (1). Assume $\bm{f^{*}}$ is not WPO. Then, $\exists\bm{f^{\prime}}$ such that ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})<\hat{\mathcal{L}}_{P}(f^{*}_{b},D_{b})\leq\epsilon_{b}\ \forall b$ (so $\bm{f^{\prime}}$ is feasible) and ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})<\hat{\mathcal{L}}_{S}(\bm{f^{*}})$. This contradicts the optimality of $\bm{f^{*}}$ among feasible $\bm{f}$; thus, $\nexists$ such $\bm{f^{\prime}}$ so $\bm{f^{*}}$ is WPO. ∎
Theorem 2.
A sequence $\bm{f^{*}}$ is PO if it is a unique solution to Problem (1) for any given $\bm{\epsilon}$.
Proof.
(Proof of Theorem 2) Let $\bm{f^{*}}$ be a unique solution to Problem (1). Then, ^$\hat{\mathcal{L}}_{S}(\bm{f^{*}})<\hat{\mathcal{L}}_{S}(\bm{f})\ \forall\bm{f}$ when ^$\hat{\mathcal{L}}_{P}(f_{b},D_{b})\leq\epsilon_{b}\ \forall b$. Assume $\bm{f^{*}}$ is not PO. Then, $\exists\bm{f^{\prime}}$ such that ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\hat{\mathcal{L}}_{P}(f^{*}_{b},D_{b})\ \forall b$, ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})\leq\hat{\mathcal{L}}_{S}(\bm{f^{*}})$, and the inequality is strict for at least one $b^{\prime}\in[B]$ or for the stability loss. If the inequality is strict for the stability loss, then we have ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})<\hat{\mathcal{L}}_{S}(\bm{f^{*}})$ (so $\bm{f^{\prime}}$ achieves a better objective value in Problem (1)) and ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\hat{\mathcal{L}}_{P}(f^{*}_{b},D_{b})\leq\epsilon_{b}\ \forall b$ (so $\bm{f^{\prime}}$ is feasible for Problem (1)). Thus, we have a contradiction with the fact that $\bm{f^{*}}$ solves Problem (1). If the inequality is strict for $b^{\prime}\in[B]$, then ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b^{\prime}},D_{b^{\prime}})<\hat{\mathcal{L}}_{P}(f^{*}_{b^{\prime}},D_{b^{\prime}})\leq\epsilon_{b^{\prime}}$, ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\hat{\mathcal{L}}_{P}(f^{*}_{b},D_{b})\leq\epsilon_{b}\ \forall b\not=b^{\prime}$, and ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})\leq\hat{\mathcal{L}}_{S}(\bm{f^{*}})$. This contradicts the uniqueness of $\bm{f^{*}}$. Thus, $\nexists$ such $\bm{f^{\prime}}$ and $\bm{f^{*}}$ is PO. ∎
In light of Theorems 1 and 2, the proposed methodology is guaranteed to produce PO sequences for models that correspond to strongly convex optimization problems (e.g., $\ell_{2}$-regularized logistic regression) and WPO sequences for all models. As we show next, to verify PO for non-strongly convex optimization problems, it suffices to solve a series of “complementary problems,” in each of which the predictive power loss for one batch is minimized subject to appropriately chosen constraints on the remaining predictive power losses and the stability loss:
Theorem 3.
A sequence $\bm{f^{*}}$ is PO if and only if it solves both Problem (1) with $\epsilon_{b}=\hat{\mathcal{L}}_{P}(f^{*}_{b},D_{b}):=\epsilon^{*}_{b}\ \forall b\in[B]$ and the following problem for all $b^{\prime}\in[B]$:
$\min_{\bm{f}}\hat{\mathcal{L}}_{P}(f_{b^{\prime}},D_{b^{\prime}})\quad\text{s.t.}\ \hat{\mathcal{L}}_{P}(f_{b},D_{b})\leq\epsilon^{*}_{b}\ \forall b\not=b^{\prime},\quad\hat{\mathcal{L}}_{S}(\bm{f})\leq\hat{\mathcal{L}}_{S}(\bm{f}^{*}).$ (2)
Proof.
(Proof of Theorem 3) Let $\bm{f^{*}}$ be a solution to Problem (1) with $\epsilon_{b}=\epsilon^{*}_{b}\ \forall b$ and the complementary Problem (2) $\forall b^{\prime}\in[B]$. Then, by optimality of $\bm{f^{*}}$ for Problem (1), $\nexists\bm{f^{\prime}}$ such that ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})<\hat{\mathcal{L}}_{S}(\bm{f}^{*})$ and ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\epsilon^{*}_{b}\ \forall b$. Further, by optimality of $\bm{f^{*}}$ for Problem (2) for all $b^{\prime}\in[B]$, $\nexists\bm{f^{\prime}}$ such that ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b^{\prime}},D_{b^{\prime}})<\epsilon^{*}_{b^{\prime}}$, ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})\leq\hat{\mathcal{L}}_{S}(\bm{f}^{*})$, and ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\epsilon^{*}_{b}\ \forall b\not=b^{\prime}$. Thus, $\bm{f^{*}}$ is PO since, in both cases, the inequality is strict for at least one objective.
For the opposite direction, let $\bm{f^{*}}$ be PO. First, assume $\bm{f^{*}}$ does not solve Problem (1) with $\epsilon_{b}=\epsilon^{*}_{b}\ \forall b$. Then, $\exists\bm{f^{\prime}}$ that is feasible, i.e., ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\epsilon^{*}_{b}\ \forall b$, such that ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})<\hat{\mathcal{L}}_{S}(\bm{f}^{*})$. This contradicts the PO of $\bm{f^{*}}$. Now, assume $\bm{f^{*}}$ does not solve Problem (2) for some $b^{\prime}$. Then, $\exists\bm{f^{\prime}}$ that is feasible, i.e., ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\epsilon^{*}_{b}\ \forall b\not=b^{\prime}$ and ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})\leq\hat{\mathcal{L}}_{S}(\bm{f}^{*})$, such that ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b^{\prime}},D_{b})<\epsilon^{*}_{b^{\prime}}$. This also contradicts the PO of $\bm{f^{*}}$. Thus, $\bm{f^{*}}$ solves the series of Problems stated in Theorem 3. ∎
In all, we can recover the entire (empirical) Pareto frontier by varying the value of the accuracy tolerance parameter in Problem (1), and possibly solving a series of complementary problems for verification. The next result, which follows by direct application of Theorem 4 from [32], guarantees that, by finding the empirical Pareto frontier, we approximately recover the true Pareto frontier with respect to all objectives:
Lemma 4.
If Assumption 1 and Observation 1 hold, then, for any $\delta\in(0,1)$, with probability at least $1-\delta$, for all PO sequences $\bm{f^{*}}$ that solve Problem (1) using the expected losses $\mathcal{L}_{P}$ and $\mathcal{L}_{S}$, there exists a PO sequence $\bm{\hat{f^{*}}}$ that solves Problem (1) using the empirical losses ^$\hat{\mathcal{L}}_{P}$ and ^$\hat{\mathcal{L}}_{S}$, such that the following hold: $\mathcal{L}_{P}(\hat{f^{*}_{b}},D_{b})\leq\mathcal{L}_{P}(f^{*}_{b},D_{b})+2C(N,\mathcal{F},\nicefrac{{\delta}}{{B}}),\ \forall b\in[B]$ and $\mathcal{L}_{S}(\bm{\hat{f^{*}}})=\mathcal{L}_{S}(\bm{f^{*}})$.
2.3 A Tractable Restricted Formulation
Depending on the type of the model used, solving Problem (1) can be challenging: the number of possible functions $f_{b}$ may grow exponentially with the number of features. This is particularly true for models that are commonly used in high-stakes healthcare applications, including decision trees and variants thereof: considering $p$ binary features and a binary classification task, then there are $2^{2^{p}}$ different ways of labeling all instances, each corresponding to a different underlying boolean function that can be represented as a decision tree. Thus, we next develop a restricted formulation that can tractably accommodate any type of ML model.
For each batch $b\in[B]$, we pre-train a finite set of candidate models, denoted as ~$\tilde{\mathcal{F}}_{b}$. Then, the problem reduces to selecting one model from each set of pre-computed models to form a sequence. Let $f_{b,j}$ be the $j$-th model in ~$\tilde{\mathcal{F}}_{b}$, and introduce binary variables $z_{b,j}$ such that $z_{b,j}=1$ if model $f_{b,j}$ is selected, and 0 otherwise. This yields the following MIO formulation:
$\begin{split}\min_{\begin{subarray}{c}f_{b,j}\in\tilde{\mathcal{F}}_{b}\\ z_{b,j}\in\{0,1\}\end{subarray}}&\quad\sum_{b=1}^{B-1}\sum_{j=1}^{|\tilde{\mathcal{F}}_{b}|}\sum_{k=1}^{|\tilde{\mathcal{F}}_{b+1}|}d(f_{b,j},f_{b+1,k})\cdot z_{b,j}\cdot z_{b+1,k}\\ \text{s.t.}&\quad\sum_{j=1}^{|\tilde{\mathcal{F}}_{b}|}\mathcal{L}_{P}(f_{b,j},D_{b})\cdot z_{b,j}\leq(1+\alpha)\cdot\min_{f\in\tilde{\mathcal{F}}_{b}}\mathcal{L}_{P}(f,D_{b})\quad\forall b\in[B]\\ &\quad\sum_{j=1}^{|\tilde{\mathcal{F}}_{b}|}z_{b,j}=1\quad\forall b\in[B].\end{split}$ (3)
The following theorem allows us to solve the restricted problem of selecting a slowly varying sequence of models (Problem (3)) in polynomial time:
Theorem 5.
Problem (3) can be reduced to a shortest path problem in a directed graph.
Proof.
(Proof of Theorem 5) Construct a directed graph $G$ as follows: Each node $v_{b,j}$ corresponds to model $f_{b,j}\in\tilde{\mathcal{F}}_{b}$ ($j$-th model for batch $b$). Include additional nodes $v_{\text{source}}$ and $v_{\text{sink}}$. Build directed edges from $v_{\text{source}}$ to $v_{1,j},\forall j,$ and from $v_{B,j}$ to $v_{\text{sink}},\forall j$, with zero weights. For $b\in[B-1]$, build directed edges from $v_{b,j}$ to $v_{b+1,k},\forall(k,j)$; assign a weight of $d(f_{b,j},f_{b+1,k})$ to each edge $(v_{b,j},v_{b+1,k})$. To enforce the accuracy tolerance constraint, for each batch $b$, include only those nodes $v_{b,j}$ corresponding to models $f_{b,j}$ that satisfy $\mathcal{L}_{P}(f_{b,j},D_{b})\leq(1+\alpha)\cdot\min_{f\in\mathcal{F}_{b}}\mathcal{L}_{P}(f,D_{b})$. Finding a sequence of models that minimizes the total stability metric while satisfying the accuracy tolerance constraint is equivalent to finding the shortest path from $v_{\text{source}}$ to $v_{\text{sink}}$ in graph $G$. Since the shortest path problem in a directed graph with non-negative weights can be solved in polynomial time (e.g., using Dijkstra’s algorithm), Problem (3) can also be solved in polynomial time. ∎
The restricted formulation (Problem (3)) is an approximation of the full formulation (Problem (1)). To assess the quality of the approximation and the impact of the size of $\mathcal{F}_{b}$ (number of candidate models per batch) on the restricted formulation, we train slowly varying sequences of linear regression models on synthetic data generated according to $y=\bm{x}^{\top}\bm{\beta}_{b}+\epsilon$, where, for each batch, the data $X\in\mathbb{R}^{n\times p}(\text{with }n=100,p=10)$ is drawn from $\bm{x}\sim\mathcal{N}(\bm{0},\Sigma),\epsilon\sim\mathcal{N}(0,\sigma)$ and the regression coefficients satisfy $|\bm{\beta}_{b}-\bm{\beta}_{0}|\leq 0.3$ for some $\bm{\beta}_{0}\in\{-1,0,1\}$. Note that, thanks to the convexity of the linear regression optimization problem, by varying the value of $\epsilon_{b}$ in Problem (1), we recover the full empirical Pareto frontier. This is visually confirmed by the solid lines in Figure 1, which reveals that the Pareto frontier is convex: with a small increase in predictive power loss, we can achieve a significant decrease in stability loss. In contrast, the restricted model does not guarantee recovery of PO solutions due to the finite sampling of models within each batch, potentially missing the PO sequence for certain accuracy tolerances. Therefore, as expected, the restricted model finds solutions that are Pareto-dominated by the ones of the full model, whereas increasing $|\mathcal{F}_{b}|$ improves the approximation. Next, we continue with a more quantitative discussion of the findings of Fig. 1.
As can be seen by the full model’s in-sample curve, we examine accuracy tolerances of $\alpha\in[0,0.01,0.03,0.1,0.3]$, where the predictive power validation loss is constrained to be within $\alpha$ of the best independent model for each batch: the corresponding gains in stability loss range between $2-12\%$. By comparing this curve, with the restricted models’ in-sample curves, we see that, for $\alpha\leq 0.03$ (which we most care about in practice), the restricted models are about $1-3\%$ worse in terms of both predictive power and stability. By comparing the in-sample with the out-of-sample curves for both the full and the restricted model, we empirically see Theorem 4’s guarantee: the in-/out-of-sample predictive power gap is around 20%, while the gaps between the full and the restricted models remain the same in-/out-of-sample.
Figure 1: Pareto frontier (in-sample and out-of-sample) between predictive power (aggregated across batches) and stability, obtained by solving Problem (1) (full) or Problem (3) (restricted for varying number of candidate models)
2.4 Model Distance Calculation
To compute the distance $d(f_{i},f_{j})$ between models $f_{i}$ and $f_{j}$, we consider two primary approaches: structural distances and feature importance-based distances.
Structural distances: This approach measures the distance by directly comparing the models’ structures. For linear/logistic regression models, we compute the distance as $d(f_{i},f_{j})=\|\bm{\beta}_{i}-\bm{\beta}_{j}\|_{2}^{2}$, where $\bm{\beta}_{i}$ and $\bm{\beta}_{j}$ are the regressions coefficient vectors of $f_{i}$ and $f_{j}$, respectively. For decision tree-based models, we use the decision tree distance metric proposed by [27] that relies on optimally matching the paths in the two trees $\mathbb{T}_{i}$ and $\mathbb{T}_{j}$, denoted as $\mathcal{P}(\mathbb{T}_{i})$ and $\mathcal{P}(\mathbb{T}_{j})$. The distance computation is as follows: first, we calculate the pairwise distance between each path in the first tree and each path in the second tree $d(p,q)$; this pairwise distance takes into account the differences in feature ranges in the two paths as well as the class labels associated with them. Then, we determine the overall distance between two decision trees by finding the optimal matching of these paths, a process that can be computed in polynomial time as a variant of the bipartite matching problem. Formally, assuming without loss of generality that $|\mathcal{P}(\mathbb{T}_{i})|\geq|\mathcal{P}(\mathbb{T}_{j})|$, and given decision variables $x_{pq}=\mathbb{I}$(path $p$ in $\mathbb{T}_{i}$ is matched with path $q$ in $\left.\mathbb{T}_{j}\right)$, the distance is defined as the following linear optimization problem:
$\min_{\begin{subarray}{c}\quad x_{pq}\in[0,1]\\ \forall(p,q)\\ \in\mathcal{P}(\mathbb{T}_{i})\times\mathcal{P}(\mathbb{T}_{j})\end{subarray}}\sum_{p\in\mathcal{P}(\mathbb{T}_{i})}\sum_{q\in\mathcal{P}(\mathbb{T}_{j})}d(p,q)x_{pq}\ \text{s.t.}\ \sum_{q\in\mathcal{P}(\mathbb{T}_{j})}x_{pq}=1,\ \forall p\in\mathcal{P}(\mathbb{T}_{i}),\ \ \sum_{p\in\mathcal{P}(\mathbb{T}_{i})}x_{pq}\geq 1,\ \forall q\in\mathcal{P}(\mathbb{T}_{j}).$ (4)
Feature importance-based distances: This approach measures the distance indirectly by comparing the models’ respective feature importances. Denoting by $\bm{\phi}_{f}\in\mathbb{R}^{p}$ the $p$-dimensional vector of feature importances that model $f$ assigns to each feature, we compute the distance as $d(f_{i},f_{j})=\|\bm{\phi}_{f_{i}}-\bm{\phi}_{f_{j}}\|_{2}^{2}$. The advantage of using feature importance-based distances is its applicability to both interpretable and black-box models. For decision tree-based models, we compute feature importances using measures intrinsic to tree growth, examining how each feature impacts node purity or classification accuracy. In particular, for CART [33, 34], we use the Gini importance, which quantifies how much a feature contributes to homogenizing the target variable within the subsets created by the splits. Given a tree $f$ with a set of nodes $V$, the Gini importance for feature $j$ is computed as $[\phi_{f}]_{j}=\frac{\sum_{v\in V}I(j,v)}{\sum_{v\in V}I(v)}$, where $I(j,v)$ represents the Gini impurity reduction at node $v$ due to splitting on feature $j$ and $I(v)$ is the total Gini impurity reduction at node $v$ regardless of the feature. For XGBoost [35], we use Gain, which focuses on the improvement in model accuracy from splits made by each feature. For general black-box models, we use SHAP (SHapley Additive exPlanations) [36], which offer a cohesive methodology to measure feature importance across different model architectures by calculating the average marginal contribution of each feature to the prediction.
2.5 Adaptive Model Retraining
We now discuss how to apply this framework in the context of adaptive model retraining: when a new batch of data becomes available and we wish to update an existing slowly varying sequence to incorporate the new data. More specifically, we have already trained an existing slowly varying sequence $f_{1},f_{2},\dots,f_{B}$ using an observational dataset $D_{B}$. We are now given an additional batch of data, which we append to the existing dataset to create $D_{B+1}$. We want to obtain a model $f_{B+1}$ that remains close in distance to the current model, $f_{B}$.
To do so, we construct an independent slowly varying sequence $f^{\prime}_{1},f^{\prime}_{2},\dots,f^{\prime}_{B+1}$ created on the updated dataset, $D_{B+1}$, and we deploy $f^{\prime}_{B+1}$. Unlike the perhaps more obvious choice of greedily choosing a new model $f^{\mathrm{greedy}}_{B+1}$ trained on $D_{B+1}$ that is close to $f_{B}$, the proposed approach achieves stability by process and generalizes to longer deployment horizons with multiple model updates and can also be used when re-training after receiving multiple data updates — e.g., constructing $f^{\prime}_{B+m}$ when the existing model is $f_{m}$. Next, we formalize the “stability by process” argument:
Definition 2.
The intra-sequence stability loss is defined as the sum of pairwise distances of adjacent models of one sequence, i.e., $\mathcal{L}_{S}(\bm{f})=\sum_{b=1}^{B-1}d(f_{b},f_{b+1})$.
Definition 3.
The inter-sequence stability loss is defined as the distance between the final models of two sequences of lengths $B$ and $B^{\prime}$ respectively, i.e., $\mathcal{L}_{S}(\bm{f},\bm{f^{\prime}})=d(f_{B},f^{\prime}_{B^{\prime}})$.
For the inter-sequence stability loss, we could alternatively compare all pairs of models in $\bm{f},\bm{f^{\prime}}$. However, such an approach (i) does not directly apply to sequences of different lengths and (ii) ignores the fact that we mostly care about the final model from each sequence (which is used in production). The next result promises that, under certain conditions, intra-sequence stability implies inter-sequence stability. In particular, we consider the implications of training a sequence of models that minimizes the intra-sequence stability loss on the inter-sequence stability when a new batch of data is introduced, and the entire sequence is retrained. We work under the additional assumption that the initial model is common in both sequences, i.e., $f_{1}=f_{1}^{\prime}=f_{0}$. This is a realistic assumption as, in practice, $f_{0}$ can be the model that was previously used in production. Further, this assumption can easily be incorporated as an extra constraint in the proposed optimization formulation.
Theorem 6.
Let $\bm{f},\bm{f^{\prime}}$ be two sequences of models trained using data $D_{1},\dots,D_{B}$ and $D_{1},\dots,D_{B},D_{B+1}$, respectively, and a stability loss metric $d:\mathcal{F}\times\mathcal{F}\mapsto\mathbb{R}_{\geq 0}$. Assume (i) their intra-sequence stability losses satisfy $\mathcal{L}_{S}(\bm{f})\leq\epsilon$ and $\mathcal{L}_{S}(\bm{f^{\prime}})\leq\epsilon^{\prime}$, and (ii) the initial models satisfy $f_{1}=f_{1}^{\prime}=f_{0}$. Then, $\mathcal{L}_{S}(\bm{f},\bm{f^{\prime}})\leq\epsilon+\epsilon^{\prime}$.
Proof.
(Proof of Theorem 6) The result follows by repeated application of the triangle inequality:
$\begin{split}\mathcal{L}_{S}(\bm{f},\bm{f\textquoteright})&:=d(f_{B},f\textquoteright_{B+1})\leq d(f_{B-1},f_{B})+d(f_{B-1},f\textquoteright_{B+1})\leq\dots\leq\sum_{b=1}^{B-1}d(f_{b},f_{b+1})+d(f_{1},f\textquoteright_{B+1})\\ &\leq\dots\leq\mathcal{L}_{S}(\bm{f})+\sum_{b=1}^{B}d(f\textquoteright_{b},f\textquoteright_{b+1})+d(f_{1},f\textquoteright_{1})\leq\mathcal{L}_{S}(\bm{f})+\mathcal{L}_{S}(\bm{f\textquoteright})+d(f_{0},f_{0})=\epsilon+\epsilon\textquoteright.\end{split}$
3 Numerical Experiments
3.1 Case Study
Dataset.
We collect daily patient data from a large hospital system in Connecticut. We extract patient features including demographics, patient status, clinical measurements, lab results, diagnoses, orders, procedures, and notes. Patient records are retrospectively identified from January 1st, 2018 to May 1st, 2022, overseeing 168,815 patients, and 865,954 patient-day records. We focus on the binary classification task of predicting patient mortality risk, defined as the likelihood of death or hospice care by hospital discharge. We evaluate model performance using the Area Under the Receiver Operating Characteristic Curve (AUC), which assesses the model’s ability to differentiate between outcomes by comparing the true positive rate against the false positive rate at various thresholds.
Data preprocessing.
In the existing production pipeline, models are trained on a 6-month basis, although the retraining window is scheduled to be reduced to 3-4 months in 2024. There are no obvious data shifts for this case study. To simulate the process of receiving data updates, we divide the 52 months into multiple sub-intervals. We follow traditional ML training procedures by dividing the dataset into training, validation, and testing sets. Specifically, all patient records admitted from September 2021 to December 2022 (16 months) are allocated as the out-of-sample testing set.
Research questions.
To assess the performance of the slowly varying methodology (SVML), we evaluate both the intra-sequence and the inter-sequence stability of the generated sequences. The former serves to illuminate the predictive power-stability trade-off that our methodology hopes to optimize, whereas the latter analyzes the stability of the adaptive model retraining strategy. Furthermore, we perform the following robustness checks: first, we investigate the sensitivity of SVML to the update frequency (or interval length); second, we examine the robustness of SVML with respect to different distance measures and the consistency of the resulting analytical insights.
Experimental methodology.
We obtain experimental results as follows. Having selected an update frequency, the number of batches $B$ is fixed. For each batch $b\in B$, we define the available training data $D_{b}$ as the first $b$ intervals and reserve the $(b+1)$-th interval as a validation set. We generate a set of pre-computed candidate models, $\mathcal{F}_{b}$, for each batch, bootstrapping to ensure diversity in the candidate models. Then, for accuracy tolerances [0.1, 0.05, 0.02, 0.01], we generate a slowly varying sequence according to the methodology outlined in Section 2.3. To benchmark the performance of these sequences, we also generate a “greedy sequence” constructed of greedily selected, best-performing models (evaluated on the validation set). We denote by $f_{b}$ the batch $b$ model chosen by the slowly varying sequence and denote by $\bar{f}_{b}$ the batch $b$ model chosen by the greedy sequence. All AUC results are reported on the 16-month testing set put aside at the start of model training.
Experimental software. We implement all algorithms in Python 3.7 and run all experiments on a standard Intel(R) Xeon(R) CPU E5-2690 @ 2.90GHz running CentOS release 7.
3.2 Results
Intra-sequence stability.
We start with an in-depth analysis of the distance-accuracy trade-off implicit in the slowly varying methodology. We select a 3-month update frequency, and conduct this experiment for three classes of models: logistic regression, CART decision trees, and XGBoost. For CART, we construct two sets of slowly varying sequences; one using a structural distance and the other using a feature importance-based distance. For each model type and each sequence, we graph the AUC and intra-sequence pairwise distance of the models comprising the slowly varying and greedy sequences in Fig. 2. Differences in test AUCs of the SVML and greedy sequences grow smaller as we reduce accuracy tolerance. Indeed, for the $\alpha=0.01$ SVML sequences, we observe only a 1.2% reduction in AUC for logistic regression and negligible reductions for the other model types. However, we note an obvious advantage in stability measured by pairwise distance, ranging from 5.5% (logistic regression) to 20% (XGBoost).
Figure 2: Test AUC (top) and distance between adjacent models (bottom) for various model/distance types
The intra-sequence stability results demonstrate the efficacy of SVML in minimizing the pairwise distance of adjacent models while sacrificing small model accuracy. By setting an accuracy tolerance of $0.01$, for example, we guarantee that the validation AUC of chosen slowly varying models is within $1\%$ of the maximum: a bound that generalizes well experimentally to the test AUC. As expected, the AUC of the models and the pairwise distances for both SVML and greedy sequences tend to, respectively, increase and decrease over time as they are given more data to be trained on.
Inter-sequence stability.
To test the inter-sequence stability of our methodology, we generate multiple different greedy and SVML XGBoost sequences (all with $\alpha=0.01)$, each using one more batch of the training data. In Table 1, we report, for both greedy and SVML, the inter-sequence stability loss between the sequences of length $b$ and $b+1$, for each $b\in[B-1]$; e.g., $T_{1-2}$ measures the inter-sequence stability loss between $(f_{1})$ and $(f_{1}^{\prime},f_{2}^{\prime})$, i.e., $d(f_{1},f_{2}^{\prime})$. SVML achieves significantly improved inter-sequence stability, which empirically verifies the stability by process argument.
Table 1: Inter-sequence stability loss for greedy and SVML sequences of various lengths
Model Name $T_{1-2}$ $T_{2-3}$ $T_{3-4}$ $T_{4-5}$ $T_{5-6}$ $T_{6-7}$ $T_{7-8}$ $T_{8-9}$ $T_{9-10}$ $T_{10-11}$
Greedy 0.476 0.455 0.405 0.363 0.400 0.397 0.353 0.377 0.284 0.255
SVML 0.561 0.418 0.317 0.340 0.294 0.323 0.328 0.233 0.284 0.210
Robustness check: update frequency.
Next, we empirically investigate the sensitivity of SVML to different update frequencies by splitting the available 52 months of data into intervals of different lengths (1, 3, 6, and 9 months). As discussed, the currently deployed 6-month frequency will change, so our experiment also serves to understand the properties of different frequencies. We choose XGBoost as the canonical model type for the experiment since this is the model in use in production, and train 25 candidate models per batch. The results, depicted in Fig. 3, demonstrate that trends and insights are consistent across the different frequencies tested. A 3-month re-training frequency exhibits more stable behavior compared to 1-month, while retaining high AUC ranging from 0.85-0.9 with improvements in distance from [0.3-0.37] to [0.275-0.21] compared to 6- or 9-month intervals.
Figure 3: AUC (top) and distance (bottom) between adjacent XGBoost models for various update frequencies.
Robustness check: distance measure.
As a final robustness check, we investigate how well the analytical insights of an SVML sequence optimized for one distance measure (XGBoost with Gain importance) generalize to a different measure (SHAP) in terms of their stability across retraining iterations. Figure 4 presents the empirical distribution of the absolute deviation in features’ SHAP values between adjacent models for both greedy and SVML sequences. Greedy achieves a mean deviation of 0.086 with a standard deviation of 0.101 and a maximum deviation of 0.625; in comparison, SVML exhibits an improved mean deviation of 0.081 with a standard deviation of 0.094 and a maximum deviation of 0.606. These results indicate that SVML’s improved stability across retraining iterations effectively translates into consistent SHAP values, affirming the method’s reliability.
Figure 4: Empirical distribution of absolute deviation in SHAP between adjacent models.
4 Conclusion
In summary, we have introduced a framework for retraining ML models, emphasizing structural stability across updates with new data. By leveraging a mixed-integer optimization approach, our methodology enables users to navigate the predictive power-stability Pareto frontier, and deploy stable models without significant compromises on predictive power. The practical benefits of our approach are highlighted in a healthcare case study, where consistent interpretability across retraining iterations is crucial for user trust and model adoption.
Limitations and future work. Depending on the underlying model type, our approach may have significant computational requirements: the bottleneck of our methodology is the formation of a sufficiently large pool of candidate models. Adaptations for scalability and efficiency might be essential for practical deployment in case studies where data and model updates are more frequent. Furthermore, the need to balance the inclusion of historical data against recent data to mitigate distributional shifts poses an additional complexity in the model retraining process. Future work will look to expand the applicability of our methods to a broader suite of models including deep learning models. The main challenges include addressing the computational needs of training such complex models, especially with the large datasets common in healthcare, and effectively measuring stability.